OpenStudy (anonymous):

The length and width of a rectangle whose length is 14cm more than it's width? PLEASE HELP!!!

OpenStudy (hockeychick23):

did they give you the total perimeter or area? thats the only way you can solve this

OpenStudy (anonymous):

No, they just gave me a rectangle and an x on the side that's all

OpenStudy (hockeychick23):

so do they just want you to write an equation for each the width and the length? because that would be: L= w+14 w=w perimeter= w+14+w+14+w+w perimeter= 4w+28

OpenStudy (hockeychick23):

thats all you can do if they don't give you the total number though, theres no way to completely solve it unless they gave you one value

OpenStudy (anonymous):

^^

OpenStudy (academicgurusinc):

I think part of the question is missing here. You cannot actually solve the entire problem but you can generate its equation. So 14x2 = 28 and you know that perimeter simply means adding up the sides. A rectangle has 4 sides and it is represented by W. so adding "W's 4 times" gets you 4w. Now all you have to do is to add everything together because the question asks for perimeter so you'll get 4w+28 as your final answer.